OpenStudy (anonymous):

Find sin θ if cot θ = - 4 and cos θ <0

OpenStudy (cwrw238):

the angle will be in the 2nd quadrant where both cot and cos are negative

OpenStudy (cwrw238):



OpenStudy (cwrw238):

so can you find sin theta?

OpenStudy (cwrw238):

sin theta = opposite / hypotenuse opposite = 1, and hypotenuse can be found by pythagoras theorem

OpenStudy (cwrw238):

do you know pythagoras theorem?

OpenStudy (cwrw238):

r u there?

OpenStudy (anonymous):

yeah sorry

OpenStudy (anonymous):

-4/1 right?

OpenStudy (cwrw238):

you can find length of hypotenuse ( longest side in right-angled triangle) using pythagoras here we have x^2 = (1)^2 + (-4)^2 where x = hypotenuse

OpenStudy (cwrw238):

no you need to find x first sin theta = 1 / x

OpenStudy (cwrw238):